OpenStudy (anonymous):

Where does the graph of y = 2x2 + 7x + 3 cross the x-axis? (−1 over 2, 0) and (−3, 0) (1 over 2, 0) and (3, 0) (−3 over 2, 0) and (−1, 0) (3 over 2, 0) and (1, 0)

OpenStudy (anonymous):

@agent0smith can you help?

OpenStudy (agent0smith):

It looks like this should factor. Try setting it equal to zero, and factoring to find where it crosses the x-axis\[\Large 2x^2 + 7x + 3 = 0\]

OpenStudy (agent0smith):

It factors like so:\[\Large (2x+1)(x+3) = 0\]

OpenStudy (anonymous):

thank you!! :)

OpenStudy (anonymous):

so I'm thinking D

OpenStudy (anonymous):

can you help me with one more??

OpenStudy (anonymous):

nevermind I got it thanks

OpenStudy (agent0smith):

\[\Large (2x+1)(x+3) = 0\]means either \[\Large 2x+1= 0\] or \[\Large x+3 = 0\]